# On independence complexes of graph products

### v1

@@ 4 Independence complexes of lexicographic products
Given two graphs $G$ and $H$, the lexicographic product $G\circ H$ is given by taking a copy of $H$ for each vertex of $G$ and adding al the possible edges between two copies if the corresponding vertices are adjacent in $G$. We now will define an analogous general construction for simplicial complexes. Given a simplicial complex $K$ with vertex set $\underline{n}$ and $\underline{L}=\{L_{1},\dots,L_{n}\}$ a family of simplicial complexes we define its polyhedral join as the simplicial complex
@@
$I(G\circ H)=\left(\underline{I(H)}\right)^{*I(G)}$
This approach to study independence complexes was used in [9, 23, 24]. In [9] it was proved that for any for any family $\underline{L}$ and a tree $T$:
@@
In [9] it was also showed that for a graph $G$ such that it is a cycle of length at least $5$, then
@@
for any any family $\underline{L}$. For a cycle of length $4$ is clear that
@@
$I(W)=\left(\underline{H^{\prime}}\right)^{*I(G^{\prime})}.$
Now, $G^{\prime}$ is a cycle or the disjoint union of paths from where we arrive to the result. ∎
@@
[9] Take $K$ a simplicial complex with vertex set $\underline{n}$ and take $(\underline{J})$ a family of simplicial complexes, then
$(\underline{J})^{*K}\simeq\bigvee_{\sigma\in K}\Sigma\mathrm{lk}(\sigma)\;\ast\bigast_{i\in\sigma}J_{i}.$
@@
$\Sigma I(W)=\Sigma\left(\underline{H^{\prime}}\right)^{*I(G^{\prime})}$
and Proposition 18 tell us that this has the homotopy type desired. ∎
Question 20.
@@ 5 Independence complexes of strong products
In this section we work on the independence complex of the strong product. We start with to results that are easy to obtain from the definition of strong product.
@@
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(u,y)]\subseteq N_{G\boxtimes H}[(v,y)]$
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(v,x)]\subseteq N_{G\boxtimes H}[(v,y)]$
Now we work on the independence complex of $P_{n}\boxtimes P_{m}$, which is known as the $n\times m$ King’s graph– given a chessboard of $n\times m$, the simplices are given by a set of kings that cannot attack each other. If $n=1$ or $m=1$, then the graph is the path. For $n=1$ and $m=2$ the graph is $K_{2}$, with the independence complex equal to $\mathbb{S}^{0}$. For $n=m=2$ the graph is $K_{4}$, therefore its independence complex is the wedge of $3$ copies of $\mathbb{S}^{0}$.
@@
$I(P_{n}\boxtimes P_{2})\simeq\Sigma I(P_{n-2}\boxtimes P_{2})\vee\bigvee_{2}\Sigma I(P_{n-3}\boxtimes P_{2})$
@@
$I(P_{n}\boxtimes P_{3})\simeq\Sigma I(P_{n-2}\boxtimes P_{3})\vee\Sigma I(P_{n-3}\boxtimes P_{3})\vee\Sigma^{2}I(P_{n-3}\boxtimes P_{3})$
@@
Now we define auxiliary graphs for the case $m=4$. For 00, we take $Q_{0}=K_{1}$ and for $n\geq 1$ we obtain $Q_{n}$ from $P_{n}\boxtimes P_{4}$ by adding a new vertex $x$ and making it adjacent to the vertices $(1,1)$ and $(1,2)$ (see Figure 1).
@@
$I(P_{n}\boxtimes P_{4})\simeq\bigvee_{3}\Sigma^{2}I(P_{n-3}\boxtimes P_{4})\vee\bigvee_{2}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4})$
$I(Q_{n})\simeq\Sigma^{2}I(P_{n-2}\boxtimes P_{4})\vee\bigvee_{3}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4})$
@@
Now we take $F^{\prime}=F-(1,3)$. We have that $N_{F^{\prime}}[(1,4)]\subseteq N_{F^{\prime}}[(2,3)]$, thus
@@
Now, in $F^{\prime}-(2,3)$ the only neighbor of $(1,4)$ is $(2,4)$, so we can erase the vertices $(3,4),(3,3)$ without changing the homotopy type– notice that the obtained graph has $2$ connected components, one which is $K_{2}$. We call this new graph $F^{\prime\prime}$. In $F^{\prime\prime}$, $N_{F^{\prime\prime}}[(1,2)]\subseteq N_{F^{\prime\prime}}[(2,2)]$, thus
@@
Using Lemma 1 is easy to obtain the following proposition.
@@
From all the work done we obtain the following corollary.
@@
Then $f_{1}(x)=1,\;f_{2}(x)=3,\;f_{3}(x)=x+2,\;g_{1}(x)=1,\;g_{2}(x)=x+2,\;g_{3}(x)=3x+1$ and for all $n\geq 4$
$f_{n}(x)=xf_{n-2}(x)+2xf_{n-3}(x)$
$g_{n}(x)=xf_{n-2}(x)+(x^{2}+x)f_{n-3}(x)$
From this is easy to get the next proposition.
@@
and for the $g_{i}^{\prime}s$ is
$G(t)=\frac{(3x+1)t^{3}+2t^{2}+t}{1-xt^{2}-(x+x^{2})t^{3}}$
@@
For $m=1$, any induced subgraph is a disjoint union of paths, therefore its independence complex is the join of independence complexes of paths, by Theorem 6 each one is contractible or has the homotopy of a single sphere.
@@
Next we take $m=3$. As before, we suppose the result is false and we take $n$ the first positive integer for which the results fails. By the work done we know $n\geq 3$. We take $H$ a vertex minimal counterexample for $n$, this means $H$ is an induced subgraph of $P_{n}\boxtimes P_{3}$ such that $I(H)$ is not contractible nor has the homotopy type of a wedge of spheres. By minimality of $n$, $H$ must have a vertex of the form $(1,j)$. If $j=1$, because $I(H)$ is not contractible at lest one of the vertices $(1,2),(2,1),(2,2)$ must be in $H$, then $N_{H}[(1,1)]$ is contained in the close neighborhood of any of this vertices. Using Lemma 1 and the minimality of $H$ we see that $H$ could not be a counterexample. Thus $j\neq 1$ and in the same way we can see that $j\neq 3$. With the same argument we can see that $(2,2)$ is not a vertex of $H$. Now, $(3,2)$neither is a vertex of $H$, if is a vertex, then $N_{H}(1,2)\subseteq N_{H}(3,2)$. With the same type of arguments we can see that $\delta(H)\geq 2$ and therefore $(2,1),(2,3),(3,1),(3,3)$ are vertices of $H$. Now $n\geq 4$, otherwise $H\cong P_{5}$ and thus $H$ is not a counterexample.
@@
    $(4,2)$ is a vertex of $H$ and $(4,1),(4,3)$ are not. In this case $H-N_{H}[(4,2)]$ the disjoin union of two graphs, one of which is isomorphic to $P_{3}$ and $H-(4,2)$ the disjoin union of two graphs, one of which is isomorphic to $P_{5}$, In this graphs the $P_{3}$ is contain in the $P_{5}$ and the other graphs are the same, thus the inclusion
[…]
@@
    $T=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).$
    Now, by Lemma 1, $I(T)\simeq I(T-(4,2))\vee\Sigma I(T-N_{T}[(4,2)])$. Where, by minimality of $H$, $I(T-N_{T}[(4,2)])$ is contractible or has the homotopy type of a wedge of spheres. We call $W=T-(4,2)$.
@@
    $W=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).$
Now, for $W$, for either of the last two cases, if $n=4$, then $I(W)\simeq*$ and $H$ is not a counterexample. Assume $n\geq 5$, then we have that
$I(W)\simeq\Sigma\left(I(W-N_{W}[(3,3)]-N_{W}[(4,3)])\cup I(W-N_{W}[(3,3)]-N_{W}[(3,1)])\right).$
Now, it is not hard to see that
$I(W-N_{W}[(3,3)]-N_{W}[(4,3)])\cup I(W-N_{W}[(3,3)]-N_{W}[(3,1)])=I(Q)$
where $Q=(W-(3,3)-(3,1)-(4,3))+(4,1)(5,3)$ if $(5,3)$ is a vertex of $W$ and if not is not a vertex of $W$ then $Q=W-(3,3)-(3,1)-(4,3)$. In either of case, $Q$ is isomorphic to a proper induced subgraph of $H$. Therefore there is no counterexample.
@@
which would imply that $H$ is not a counterexample. Therefore $(1,1)$ is not a vertex of $H$. Similarly $(1,4)$ is not a vertex of $H$. Now, at least one of $(1,2)$ and $(1,3)$ is in $H$, otherwise $H$ would be an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. First assume both vertices are in $V(H)$, then $(2,2)$ and $(2,3)$ are not in $V(H)$. Both $(1,2)$ and $(1,3)$ must have degree $2$, otherwise $I(H)\cong\Sigma I(H-N_{H}[(1,2)])$ or $I(H)\cong\Sigma I(H-N_{H}[(1,3)])$ and $H$ would not be a counterexample. Then $H$ is like Figure 2(a), where we know that the solid vertices and edges are in $H$ and the other may be. By Theorem 4 $I(H)\simeq\Sigma I(W)$ where $W$ is the graph in Figure 2(b) obtained from $H-(1,2)-(1,3)-(2,4)$ by adding the edges $\{(2,1),(3,b)\}$ if $(3,b)$ is in $H$. If $(3,1)$ or $(3,4)$ are not in $W$, the $W$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. Then $(3,1)$ and $(3,4)$ are in $W$.
If $(3,2)$ is in $V(W)$, then $I(W)\simeq I(W-(3,2))\vee\Sigma I(W-N_{W}[(3,2)])$. Now, $W-N_{W}[(3,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$, thus we only are interested in $W-(3,2)$. If $(3,3)$ is in $W$, then $I(W-(3,3))\simeq I(W-(3,2)-(3,3))\vee\Sigma I(W-(3,2)-N_{W}[(3,3)])$ and again, we are only interested in $W-(3,2)-(3,3)$. So we can assume that neither $(3,2)$ nor $(3,3)$ are in $W$. If both $(4,1)$ and $(4,2)$ are in $W$, then $I(W)\simeq I(W-(4,2))\vee\Sigma I(W-N_{W}[(4,2)])$, where as before we are only interested in $W-(4,2)$. Thus we can assume that only one of $(4,1)$ and $(4,2)$ is in $W$. By the same argument we can assume that only one of $(4,3)$ and $(4,4)$ is in $W$. Up to isomorphism there are only three possible cases shown in Figures 2(c), 2(e) and 2(e).
@@
Case (ii): By Theorem 5, $I(W)\simeq\Sigma I(R)$ where $R$ is obtained from $W-N_{W}[(2,1)]$ by adding the edge $\{(4,1),(4,4)\}$ (see Figure 3(a)). With arguments similar at what we have done, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $R$, the same for $(5,3)$ and $(5,4)$. Therefore there are four possible cases shown in Figures 3(b), 3(c) and 3(d). Case (a): This case is not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. Case (b): In this case we have that $N_{R}((4,1))\subseteq N_{R}((5,3))$ and $N_{R}((4,4))\subseteq N_{R}((5,2))$, thus $I(R)\simeq I(R-(5,2)-(5,3))$. This can not happen as $R-(5,2)-(5,3)$ is the disjoint union of an edge and an induced subgraph of a graph isomorphic to $P_{n-5}\boxtimes P_{4}$. Case (c): With the same arguments as before we can assume that only one vertex of $(6,1)$ and $(6,2)$ is in $R$, the same for $(6,3)$ and $(6,4)$. Thus there are four cases shown in Figures 3(e), 3(f), 3(g) and 3(h). The first two cases are not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. In other two cases, by Theorem 4, $I(R)\simeq\Sigma I(R-(4,1)-(4,4)-(5,1))$, but $R-(4,1)-(4,4)-(5,1)$ is an induced subgraph of a graph isomorphic to $P_{n-4}\boxtimes P_{4}$. From all that we get that case (ii) is not possible.
@@
Case (iii): As we have done before, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $R$, the same for $(5,3)$ and $(5,4)$. There are four possible cases shown in Figures 4(a), 4(b), 4(c) and 4(d). In any case, by Theorem 5, we can add the edge $\{(4,1),(4,3)\}$ to obtain a graph $T$ such that $I(W)\simeq\Sigma I(T)$. In the first case we have that $N_{T}((4,1))\subseteq N_{T}((5,3))$, thus $I(H)\simeq I(T-(5,3))$ by Lemma 1. But $T-(5,3)$ is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefore this case is not possible. In the next two case we have that $T$ is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefores these case are not possible. If $n=5$, then $T\cong P_{4}$ and $I(T)\simeq*$. Then $n\geq 6$. By Theorem 4, $I(T)\simeq\Sigma I(W^{\prime})$, with $W^{\prime}$a graph like the graph in Figure 2(b) and we are again in case (iii). If $n=6$, then $W^{\prime}\cong P_{5}$ (see Figure 2(e)) and $I(W^{\prime})\simeq\mathbb{S}^{1}$. If $n=7$, then $W^{\prime}\cong P_{7}$ and $I(W^{\prime})\simeq*$. If $n=8$ using Theorem 4 we get a graph similar to the graph in Figure 2(b). We can keep doing this, but the graph is finite, thus we see that this case is not possible.
@@
Lastly, we assume that $(1,2)$ is in $V(H)$ and $(1,3)$ is not. If $(2,2)$ is a vertex of $H$, by Lemma 1 we have that
@@
and we are only interested in $H-(2,2)$. So we can assume that $(2,2)$ is not a vertex of $H$. Then $(2,1)$ and $(2,3)$ are in $V(H)$, otherwise $(1,2)$ would have degree $1$ or 00 and $H$ would not be a counterexample. If $(2,4)$ is a vertex of $H$, then
@@
where we are only interested in $I(F-(4,2))$. Thus we can assume that $(4,2)$ is not a vertex of $F$. There are two possible cases. If $(4,4)$ is in $V(F)$, then $F$ would be an induced subgraph of a graph isomorphic to $P_{n-2}\boxtimes P_{4}$. Assume that $(4,3)$ is in $V(F)$. If $n=4$, then $F\cong P_{4}$ and $I(F)\simeq*$. Thus $n\geq 5$. By Theorem 4 $I(F)\simeq\Sigma I(F^{\prime})$ where $F^{\prime}$ is like the graph in Figure 2(b) and we are in a case already done. Therefore there is no counter example for the result. ∎
Question 30.
@@
$I(G-N_{G}[(1,2)])\simeq\Sigma I(G-N_{G}[(1,2)]-(3,2))\vee\Sigma I(G-N_{G}[(1,2)]-N_{G}[(3,2)])$
@@
    For $1$, $s_{1}$ tell us if we get the complex form $I(G-(1,2))$ or $I(G-N_{G}[(1,2)])$.
@@
    For $i\geq 2$, $s_{i}$ tell us if in the step $i$ we removed $(s_{1}+\cdots+s_{i-1}+1,2)$ or its closed neighborhood.

### v2

@@ 4 Independence complexes of lexicographic products
Given two graphs $G$ and $H$, the lexicographic product $G\circ H$ is given by taking a copy of $H$ for each vertex of $G$ and adding all the possible edges between two copies if the corresponding vertices are adjacent in $G$. We now will define an analogous general construction for simplicial complexes. Given a simplicial complex $K$ with vertex set $\underline{n}$ and $\underline{L}=\{L_{1},\dots,L_{n}\}$ a family of simplicial complexes we define its polyhedral join as the simplicial complex
@@
$I(G\circ H)=\left(\underline{I(H)}\right)^{*I(G)}.$
This approach to study independence complexes was used in [poljoin, MR4477852, okurajoinbosq]. In [poljoin] it was proved that for any for any family $\underline{L}$ and a tree $T$:
@@
In [poljoin] it was also showed that for a graph $G$ such that it is a cycle of length at least $5$, then
@@
for any any family $\underline{L}$. For a cycle of length $4$ it is clear that
@@
$I(W)=\left(\underline{I(H^{\prime})}\right)^{*I(G^{\prime})}.$
Now $G^{\prime}$ is a cycle or a forest, thus $I(W)$ is contractible or it has the homotopy type of a wedge of spheres or it has the homotopy type of a disjoint union of wedges of spheres. ∎
@@
[poljoin] If $K$ is a simplicial complex with vertex set $\underline{n}$ and $(\underline{J})$ is a family of simplicial complexes, then
$\Sigma(\underline{J})^{*K}\simeq\bigvee_{\sigma\in K}\Sigma\mathrm{lk}(\sigma)\;\ast\bigast_{i\in\sigma}J_{i}.$
@@
$\Sigma I(W)=\Sigma\left(\underline{H^{\prime}}\right)^{*I(G^{\prime})}.$
By Proposition 20, $\Sigma I(W)$ has the desired homotopy type. ∎
Question 22.
@@ 5 Independence complexes of strong products
In this section we focus on the independence complex of strong products of paths. We start with two results that are easy to obtain from the definition of strong product.
@@
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(u,y)]\subseteq N_{G\boxtimes H}[(v,y)],$
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(v,x)]\subseteq N_{G\boxtimes H}[(v,y)].$
The graph $P_{n}\boxtimes P_{m}$ is known as the $n\times m$ King’s graph– given a chessboard of $n\times m$, an independent sets is given by a set of kings that cannot attack each other. If $n=1$ or $m=1$, then the graph is a path. For $n=1$ and $m=2$ the graph is $K_{2}$, with the independence complex equal to $\mathbb{S}^{0}$. For $n=m=2$ the graph is $K_{4}$, therefore its independence complex is the wedge of $3$ copies of $\mathbb{S}^{0}$.
@@
$I(P_{n}\boxtimes P_{2})\simeq\Sigma I(P_{n-2}\boxtimes P_{2})\vee\bigvee_{2}\Sigma I(P_{n-3}\boxtimes P_{2}).$
@@
$I(P_{n}\boxtimes P_{3})\simeq\Sigma I(P_{n-2}\boxtimes P_{3})\vee\Sigma I(P_{n-3}\boxtimes P_{3})\vee\Sigma^{2}I(P_{n-3}\boxtimes P_{3}).$
@@
Now we define auxiliary graphs for the case $m=4$. For 0, we take $Q_{0}=K_{1}$ and for $n\geq 1$ we obtain $Q_{n}$ from $P_{n}\boxtimes P_{4}$ by adding a new vertex $x$ and making it adjacent to the vertices $(1,1)$ and $(1,2)$ (see Figure 1).
@@
$I(P_{n}\boxtimes P_{4})\simeq\bigvee_{3}\Sigma^{2}I(P_{n-3}\boxtimes P_{4})\vee\bigvee_{2}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4}),$
$I(Q_{n})\simeq\Sigma^{2}I(P_{n-2}\boxtimes P_{4})\vee\bigvee_{3}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4}).$
@@
We take $F^{\prime}=F-(1,3)$. We have that $N_{F^{\prime}}[(1,4)]\subseteq N_{F^{\prime}}[(2,3)]$, thus
@@
Now, in $F^{\prime}-(2,3)$ the only neighbor of $(1,4)$ is $(2,4)$, so we can erase the vertices $(3,4),(3,3)$ without changing the homotopy type– observe that the graph obtained has $2$ connected components, one of these components is isomorphic to $K_{2}$. We call this new graph $F^{\prime\prime}$. In $F^{\prime\prime}$, $N_{F^{\prime\prime}}[(1,2)]\subseteq N_{F^{\prime\prime}}[(2,2)]$, thus
@@
Using Lemma 1 it is easy to obtain the following proposition.
@@
where $\tilde{\beta}_{i}$ is the reduced $i$-Betti number. Then $f_{1}(x)=1,\;f_{2}(x)=3,\;f_{3}(x)=x+2,\;g_{1}(x)=1,\;g_{2}(x)=x+2,\;g_{3}(x)=3x+1$ and for all $n\geq 4$
$f_{n}(x)=xf_{n-2}(x)+2xf_{n-3}(x),$
$g_{n}(x)=xf_{n-2}(x)+(x^{2}+x)f_{n-3}(x).$
From this it is easy to get the following proposition.
@@
and for the $g_{i}^{\prime}s$ the generating function is
$G(t)=\frac{(3x+1)t^{3}+2t^{2}+t}{1-xt^{2}-(x+x^{2})t^{3}}.$
Now we give the analogous theorem to Theorem 17 for the strong product.
@@
For $m=1$, any induced subgraph is a disjoint union of paths, therefore its independence complex is the join of independence complexes of paths, by Theorem 6 each one is contractible or has the homotopy type of a single sphere.
@@
Next we take $m=3$. As before, we suppose the result is false and we take $n$ the first positive integer for which the results fails. By the work done we know $n\geq 3$. We take $H$ a vertex minimal counterexample for $n$, this means $H$ is an induced subgraph of $P_{n}\boxtimes P_{3}$ such that $I(H)$ is not contractible nor has the homotopy type of a wedge of spheres. By minimality of $n$, $H$ must have a vertex of the form $(1,j)$. If $j=1$, because $I(H)$ is not contractible at least one of the vertices $(1,2),(2,1),(2,2)$ must be in $H$, then $N_{H}[(1,1)]$ is contained in the close neighborhood of any of these vertices. Using Lemma 1 and the minimality of $H$ we see that $H$ could not be a counterexample. Thus $j\neq 1$. With a similar argument we can see that $j\neq 3$. Then $(1,2)$ is a vertex of $H$. Now $(2,2)$ is not a vertex of $H$, otherwise $N_{H}[(1,2)]\subseteq N_{H}[(2,2)]$. Now $(3,2)$ is not a vertex of $H$, because if $(3,2)$ is a vertex of $H$, then $N_{H}(1,2)\subseteq N_{H}(3,2)$ and $I(H)\simeq I(H-(3,2))$. With the same type of arguments we can see that $\delta(H)\geq 2$ and therefore $(2,1),(2,3),(3,1),(3,3)$ are vertices of $H$. Now $n\geq 4$, otherwise $H\cong P_{5}$ and $H$ would not be a counterexample.
@@
    $(4,2)$ is a vertex of $H$ and $(4,1),(4,3)$ are not. In this case $H-N_{H}[(4,2)]$ the disjoin union of two graphs, one of which is isomorphic to $P_{3}$ and H(4,2)H-(4,2) the disjoin union of two graphs, one of which is isomorphic to $P_{5}$, In this graphs the $P_{3}$ is contain in the $P_{5}$ and the other graphs are the same, thus the inclusion
    $I(H-N_{H}[(4,2)])\lhook\mathrel{\mkern-3.0mu}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow I(H-(4,2))$
@@
    T=(H(1,2)(2,1)(2,3))+(3,1)(3,3).T=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).
    Now, by Lemma 1, $I(T)\simeq I(T-(4,2))\vee\Sigma I(T-N_{T}[(4,2)])$. Observe that $T-N_{T}[(4,2)]$ is an induced subgraph of $H$, therefore, by the minimality of $H$, $I(T-N_{T}[(4,2)])$ is contractible or has the homotopy type of a wedge of spheres. We take W=T(4,2)W=T-(4,2).
@@
    W=(H(1,2)(2,1)(2,3))+(3,1)(3,3).W=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).
For either of the last two cases, if $n=4$, then $I(W)\simeq*$ and $H$ is not a counterexample. Assume $n\geq 5$, by Theorem 4 we have that $I(W)\simeq\Sigma I(Q)$ where Q=(W(3,3)(3,1)(4,3))+(4,1)(5,3)Q=(W-(3,3)-(3,1)-(4,3))+(4,1)(5,3) if $(5,3)$ is a vertex of $W$ and if not is not a vertex of $W$ then Q=W(3,3)(3,1)(4,3)Q=W-(3,3)-(3,1)-(4,3). In either of case, $Q$ is isomorphic to a proper induced subgraph of $H$. Therefore there is no counterexample.
@@
which would imply that $H$ is not a counterexample. Therefore $(1,1)$ is not a vertex of $H$. Similarly $(1,4)$ is not a vertex of $H$. Now, at least one of $(1,2)$ and $(1,3)$ is in $H$, otherwise $H$ would be an induced subgraph of $P_{n+1}\boxtimes P_{4}-(1,1)-(1,2)-(1,3)-(1,4)\cong P_{n}\boxtimes P_{4}$. First assume both vertices are in $V(H)$, then $(2,2)$ and $(2,3)$ are not in $V(H)$. Both $(1,2)$ and $(1,3)$ must have degree $2$, otherwise $I(H)\cong\Sigma I(H-N_{H}[(1,2)])$ or $I(H)\cong\Sigma I(H-N_{H}[(1,3)])$ and $H$ would not be a counterexample. Then $H$ is like Figure 2(a), where we know that the solid vertices and edges are in $H$ and the other may be– this will be the convention for all the figures. By Theorem 4 $I(H)\simeq\Sigma I(W)$ where $W$ is the graph in Figure 2(b) obtained from H(1,2)(1,3)(2,4)H-(1,2)-(1,3)-(2,4) by adding the edges $\{(2,1),(3,b)\}$ if $(3,b)$ is in $H$. If $(3,1)$ or $(3,4)$ are not in $W$, the $W$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. Then $(3,1)$ and $(3,4)$ are in $W$.
If $(3,2)$ is in $V(W)$, then $I(W)\simeq I(W-(3,2))\vee\Sigma I(W-N_{W}[(3,2)])$. Now, $W-N_{W}[(3,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$, thus we only are interested in W(3,2)W-(3,2). If $(3,3)$ is in $W$, then $I(W-(3,3))\simeq I(W-(3,2)-(3,3))\vee\Sigma I(W-(3,2)-N_{W}[(3,3)])$ and again, we are only interested in W(3,2)(3,3)W-(3,2)-(3,3). So we can assume that neither $(3,2)$ nor $(3,3)$ are in $W$. If both $(4,1)$ and $(4,2)$ are in $W$, then $I(W)\simeq I(W-(4,2))\vee\Sigma I(W-N_{W}[(4,2)])$, where as before we are only interested in W(4,2)W-(4,2). Thus we can assume that only one of $(4,1)$ and $(4,2)$ is in $W$. By the same argument we can assume that only one of $(4,3)$ and $(4,4)$ is in $W$. Up to isomorphism there are only three possible cases shown in Figures 2(c), 2(d) and 2(e).
@@
Case (ii): By Theorem 5, $I(W)\simeq\Sigma I(R)$ where $R$ is obtained from $W-N_{W}[(2,1)]$ by adding the edge $\{(4,1),(4,4)\}$ (see Figure 3(a)). With arguments similar at what we have done, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $R$, the same for $(5,3)$ and $(5,4)$. Therefore there are four possible cases shown in Figures 3(b), 3(c) and 3(d). Case (a): This case is not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. Case (b): In this case we have that $N_{R}((4,1))\subseteq N_{R}((5,3))$ and $N_{R}((4,4))\subseteq N_{R}((5,2))$, thus $I(R)\simeq I(R-(5,2)-(5,3))$. This can not happen as R(5,2)(5,3)R-(5,2)-(5,3) is the disjoint union of an edge and an induced subgraph of a graph isomorphic to $P_{n-5}\boxtimes P_{4}$. Case (c): With the same arguments as before we can assume that only one vertex of $(6,1)$ and $(6,2)$ is in $R$, the same for $(6,3)$ and $(6,4)$. Thus there are four cases shown in Figures 3(e), 3(f), 3(g) and 3(h). The first two cases are not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. In other two cases, by Theorem 4, $I(R)\simeq\Sigma I(R-(4,1)-(4,4)-(5,1))$, but R(4,1)(4,4)(5,1)R-(4,1)-(4,4)-(5,1) is an induced subgraph of a graph isomorphic to $P_{n-4}\boxtimes P_{4}$. From all that we get that case (ii) is not possible.
@@
Case (iii): As we have done before, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $W$, the same for $(5,3)$ and $(5,4)$. There are four possible cases shown in Figures 4(a), 4(b), 4(c) and 4(d). In any case, by Theorem 5, we can add the edge $\{(4,1),(4,3)\}$ to obtain a graph $T$ such that $I(W)\simeq\Sigma I(T)$. In the first case we have that $N_{T}((4,1))\subseteq N_{T}((5,3))$, thus $I(H)\simeq I(T-(5,3))$ by Lemma 1. But T(5,3)T-(5,3) is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefore this case is not possible. In the next two case we have that $T$ is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefore these cases are not possible. Thus the only possible case is Figure 4(d). If $n=5$, then $T\cong P_{4}$ and $I(T)\simeq*$. Then $n\geq 6$. By Theorem 4, $I(T)\simeq\Sigma I(W^{\prime})$, with $W^{\prime}$a graph like the graph in Figure 2(b). If $n=6$, then $I(W^{\prime})$ is homotopy equivalent to $\mathbb{S}^{0}$ or to $\mathbb{S}^{0}\vee\mathbb{S}^{0}$. Thus $n\geq 7$. We can proceed as at the beginning of case (iii) until we are again in the case of Figure 4(d). If $n=7$, then $W^{\prime}\cong P_{5}$ and $I(W^{\prime})\simeq\mathbb{S}^{1}$.Therefore $n\geq 8$ and, as before, by Theorems 5 and 4 we get a graph similar to the graph in Figure 2(b). We can keep doing this, but the graph is finite, thus this case is not possible.
@@
Therefore only one of $(1,2)$ and $(1,3)$ is in $H$. We assume that $(1,2)$ is in $V(H)$ and $(1,3)$ is not. If $(2,2)$ is a vertex of $H$, by Lemma 1 we have that
@@
and we are only interested in H(2,2)H-(2,2). So we can assume that $(2,2)$ is not a vertex of $H$. Then $(2,1)$ and $(2,3)$ are in $V(H)$, otherwise $(1,2)$ would have degree $1$ or 0 and $H$ would not be a counterexample. If $(2,4)$ is a vertex of $H$, then
@@
where we are only interested in I(F(4,2))I(F-(4,2)). Thus we can assume that $(4,2)$ is not a vertex of $F$. There are two possible cases. If $(4,4)$ is in $V(F)$, then $F$ would be an induced subgraph of a graph isomorphic to $P_{n-2}\boxtimes P_{4}$. Assume that $(4,3)$ is in $V(F)$. If $n=4$, then $F\cong P_{4}$ and $I(F)\simeq*$. Thus $n\geq 5$. By Theorem 4 $I(F)\simeq\Sigma I(F^{\prime})$ where $F^{\prime}$ is like the graph in Figure 2(b). Therefore there is no counter example. ∎
Question 32.
@@
and
$I(G-N_{G}[(1,2)])\simeq\Sigma I(G-N_{G}[(1,2)]-(3,2))\vee\Sigma I(G-N_{G}[(1,2)]-N_{G}[(3,2)]).$
@@
    For $1$, $s_{1}$ tell us if we get the complex form I(G(1,2))I(G-(1,2)) or $I(G-N_{G}[(1,2)])$.
@@
    For $i\geq 2$, $s_{i}$ tell us if in the step $i$ we erased $(s_{1}+\cdots+s_{i-1}+1,2)$ or its closed neighborhood.